NAOJ GW Elog Logbook 3.2

I checked the gluing test with the uncoated half ball (elog entry http://www2.nao.ac.jp/~gw-elog/osl/?r=390).
I pushed a bit the half ball and I couldn't remove it, so the gluing is successful.
Therefore I could proceed with the gluing of the coated one.
I made a hole in a plastic cap to hold the coated half ball.
Put some glue on the border oh the central hole of a 1" metal plate
Took the glued plate and attached on the half ball.
Used the plastic cap to adjust the position.
Now I wait for the glue to dry and to cure.

PICOMOTORS CONTROL
Today we completed the test on picomotor motion succeding in moving a test picomotor using "old" picomotor Multiaxis driver 8732 (see picture1) controlled by pc.
A recap of the different option we tested to move the picomotrs:
1) by using "new" newfocus 8753 picomotor controller, connected to the pc by ethernet connection and controlled using labview (MCL commands). See picture 2.
2) by using Multiaxis driver 8732 controlled by hand-pad.
3) by using Multiaxis driver 8732 conected to the computer by GPIB (using GPIB-USB -HS cable) and custom labview VI.
PICOMOTORS CABLING

But if everything works well, the Faraday Isolator should stop the reflect beams. So we checked the polarization again, try to get pure p or s polarization at the place we want. After that, the situation gets better, the peak looks normal, and the green stopped flashing. But the error signal sometimes still have some high frequency oscillation.(To further check tomorrow)
Before the two steering mirror, one of them is dichoric mirror and the other is normal mirror, for get rid of the green reflection, we changed the normal mirror with a second dichoric mirror.
Then we check the power everywhere, we have about 490mW at the output of the laser, 350 after the EOM and 230mW enter the cavity. So now nearly half of the power reach the cavity. With not very good alignment, we got 95mW in green finally, so now the green efficiency is 35%-40%. We try to measure the power reflect back after one of the dichoric mirror, 5% of green power is reflected.

I have 3 sets of roughness measurement data from LMA. It is a sample of silica substrate used to transfer crystalline coating on it. The sets are taken at different points and surfaces of the same sample. I plot the measurement after tilt and curvature removal. I plot the PSDs of the roughness. Result is that the roughness is uniform on the sample. RMS=1nm

Participants: Eleonora, Yuefan, Manuel, Matteo Tacca, Raffaele, Tatsumi-san
Today Tatsumi-san gave us a practice class about how to glue magnets and stand-off ( a.k.a wire braker or spacers) to the mirrors.
Pictures showing the procedure can be found in the following link
https://www.dropbox.com/sh/ggpodec805a30k3/AAB0Bd2UBSd00-coYK7Wp5QYa?dl=0
In the attached file you can find the instruction to make the glue. The glue can be used up to 1 hour after its preparation and the glued parts has to wait for at least 24 hours.
The procedure for Kagra mirror (which is not so different) can be found here
http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=4248

There are some problems we need to fixed,
1. The crystal is not fixed, when we try to get the mode matching, sometimes cannot avoid touch the cavity, then we lost everything. Matteo did the alignment many times in few hours working.
2.There are too many green reflected back, the maximum efficiency we got is about 17%, with a input infrared power of 670mW, and output green of 119mW, when the cavity is locked. But there are at least four reflect green beam on the mirror before the dichoric mirror, we are able to measure two of them, one has 1-2 mW power, the other has 4mW.
3. The green beam has some fringe on it, so this maybe means there is not only one green beam come out from the cavity, but we are not sure.
After first time we got a efficiency of about 16%, we break the lock and wait for twenty minutes to see if the system is stable or not. When we try to lock it again 20min later, the output of green is almost the same with some tiny re-alignment.

I finished the configuration of the optical setup for our hemispherical scatterometer Yesterday.
The beam of the new laser is now focussed to a beam diameter of ca. 0.32 mm at the sample's surface.
However, I think there are still things that need to be done:
A diagram of measurements of the beam's profile is shown in the attachment. It shows the measured values of the profile without a focussing lens (circles) and with a lens at 430 mm distance to the collimator (stars). The actual distance of the lens in the setup is, however, 490 mm. Due to a small mistake, I underestimated the distance to the sample a little bit and had to readjust the positionof the lens.
Anyway, as can be seen, the measured values are in consistence with the calculated ones (dotted curves) only until the waist has been reached. With increasing distance again, there is a considerable difference between measured and theoretical value of the beam's radius*.
I think I have to measure again and try to find out the reason of the difference.
*The values are related to a Gaussian fit of the beam's cross-sectional intensity and represent the diameter of a Gaussian curve.

https://www.thorlabschina.cn/newgrouppage9.cfm?objectgroup_id=4808&pn=BSX17
And we test the cable which can send signal from end room to the central room, we followed the cable to the end room, found a same box as in the central room, send a sine signal from there and got the signal with the oscilloscope in the central.But in the west end, we could not find the box, but I think we will not use that arm.

On Friday, Manuel and Takahashi-san went to TAMA and sucessfully tested picomotor controller connection. Here a summary of what I have understood about this:
In order to make the computer be able to comunicate with the picomotor controller they must have the same IP address (except for the last digits).
Controller should have a static iP address and for this reason cannot be connected directly to NAOJ network (since it doesn't allow a static IP address). We have to choose a port in the getaway which are isolated from NAOJ network (and disconnetected from the internet).
The network used for the optical lever control is one if the and can be a good choise. The relative IP adreddes are of the form 133.140.121.XX.
The ports relative to this network in the gataway that can be found up-left corner in tama central room are the number 1 and 2 (see pictures 1 and 2).
The steps to establish the connection :
1) Pick a gateaway port which is isolated from the naoj network (disconnected from the internet) and find the relative IP addreses
2) Using serial port assign to the controller a proper IP address ( command IPADRR
3) The proper IP address has to be the same of the network except for the last digits which has to be not yet assigned ( make a ping to check..)
5) Also the pc IP adress should be check and eventually changed (with the " assign manually option") to a proper one. (Same of the network with different last digits)
6) Connect both the PC and the controller to the same network and try the ping
[NB if we are not interested in being connected to a network we can simply connect the pc and the controller to the same switch and manually assign them a compatible IP address. In order to connect the PC directy to the controller a crossover cable is needed ]

following Takahashi-san's advice I connected PC, controller and network cable to the same Ethernet hub. The ping still fails. BUT if I unplug the nework cable, which means I isolate my system from NAOJ network, then I get this ping returns:
Pinging 10.68.10.241 with 32 bytes of data:
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Ping statistics for 10.68.10.241:
Packets: Sent = 4, Recieved = 4, Lost = 0 (0% loss),
In the case of 192.168.11.121-IP controller the ping still fails, even in the isolated case.

I found that NAOJ network uses DNS, so I follow the instructions of "Ethernet Configuration Example 2" at page 50 of the manual. (last time we tried Example 1 and, apparently, it didn't work because of DNS).
Through serial port, the IPADDR command returns 192.168.11.121 for one controller and 10.68.10.241 for the controller labeled NM2.
I connect the controller with IP address 192.168.11.121.
HOSTNAME command returns: nf8752-295333
I change the hostname with the command HOSTNAME NF8752-001
Save the configuration with the SAV command
RESET
Try to ping NF8752-001, failed (it says: "Ping request could not find host NF8752-001")

In the second figure, the y-axis is the green power over infrared power, I just want to show the efficiency of infrared change into green at different input power.
I already changed the dichoric mirror from Thorlab two inch one into CVI one inch mirror has a anti-reflect coating on the back side can reduce the amount of power reflect by it. Still doing the alignment with the new mirror.

In the past days we made some tests in order to check the communication between the computer and the picomotor controller.
Details on the picomotors drivers can be found in the user's manual at this link:
http://164.54.212.3/control_systems_manuals/87XX_Manual_RevC.pdf
We followed the instructions to configure the ethernet connection at page 48/50 of the manual but we didn't succeed in setting the communication.
We did the following steps:
1) Connect a PC to the serial port and send a IPADDR MCL command to get the IP address of the controller. For the controller named NM2 we got the IP 10.68.10.241.
2) Connect the controller and the computer to the same ethernet network in TAMA and try to ping the drivers from the pc without success.
3) Since we are convinced that the IP we got is not the good one to connect to TAMA newtwork ( all the other adresses are of the form 133.40.XXX.XX) we tried to ask the IP adress (again by using serial port) after having connected the drivers to the ethernet but we got the same answer.
We probably need the advice of a TAMA network expert in order to fix the problem.
I found that NAOJ network uses DNS, so I follow the instructions of "Ethernet Configuration Example 2" at page 50 of the manual. (last time we tried Example 1 and, apparently, it didn't work because of DNS).
Through serial port, the IPADDR command returns 192.168.11.121 for one controller and 10.68.10.241 for the controller labeled NM2.
I connect the controller with IP address 192.168.11.121.
HOSTNAME command returns: nf8752-295333
I change the hostname with the command HOSTNAME NF8752-001
Save the configuration with the SAV command
RESET
Try to ping NF8752-001, failed (it says: "Ping request could not find host NF8752-001")
following Takahashi-san's advice I connected PC, controller and network cable to the same Ethernet hub. The ping still fails. BUT if I unplug the nework cable, which means I isolate my system from NAOJ network, then I get this ping returns:
Pinging 10.68.10.241 with 32 bytes of data:
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Reply from 133.40.117.65: Destination host unreachable.
Ping statistics for 10.68.10.241:
Packets: Sent = 4, Recieved = 4, Lost = 0 (0% loss),
In the case of 192.168.11.121-IP controller the ping still fails, even in the isolated case.

We have 2 crystalline coating samples from CMS company, one attached on a silica substrate and one on a sapphire substrate.
I plot the surface roughness of the crystalline coating samples we measured at LMA.
I calculated the PSD of the surface profile.

Then I put the EOM back and adjust it to a good position, also change the move the BS mirror a little bit for not let the beam goes too closed to the edge. After these, I do the alignment again and get pretty good mode matching.
I changed the laser power, to see with different amount of infrared what is the efficiency of the green production. And at one infrared power I measured the green beam with different SHG temperature, and found the best phase matching temperature. I tried to make a plot, but found at there are two points are not very reasonable, I will measure it again tomorrow to see if the problem comes from the measurement.

Then the EOM is very sensitive, even touch it can change the transmit power. There are some possibility, one is while touching, the alignment changed, the other one is there are some electric effect, but not quiet sure about this guess. Or it maybe caused by the polarization before the EOM.
So I will check the following things: the alignment of the EOM, the polarization between the laser and the EOM.
Checking the alignment of EOM is also because we lose too much power there, and also between the waveplate, I have already check the waveplate all have AR coating. At the very beginning of the experiment, I checked the transmission rate of the Faraday Isolator and EOM are both less than 90%.
There are also a small problem with the BS mirror, the beam is near the edge of the mirror which will lead to more scattering. So I will also adjust the position of it.

The 300m arm tube is in vacuum, 71mTorr at the South End and 92mTorr at the Mid point.
Three TMPs (Near, Mid, and End) along the arm can work with the rotary pumps (RPs).
We can start to evacuate the tube without initial pumping by a large RP.
The next step;
1) Start evacuation.
2) Check the penning gauges in HV (<10^-4 Torr).
3) Check the ionization pumps (IPs) in UHV (<10^-7 Torr).


before gluing the coated spherical mirrors to the 1"-support, I made a test with an uncoated one. It is cured with UV light, so I put it under a neon light wich have a small part of UV. The higher the UV intensity, the faster the cure, so now I have to wait some days for curing and bonding.

Then I checked the infrared power at different position with less output power of laser.The result:
Laser-(808.3mW)-waveplate-(795.5)-waveplate-(794)-Faraday Isolator-(728.4)-waveplate-(650)-EOM-(517)-BS mirror-\(429.4)-(399,6)-lens1-(331.9)
-mirror/-(323.2)-lens2-(307.9)-waveplate-(303.0)-lens3-(299.2)-mirror\-(293.7)-dichoric mirror-/-(292.8)-SHG
Between the BS mirror and the lens1, the power seems reduce with the distance.
And after checked the power, I found out I lost the alignment seriously. No matter how I change, the TEM00 cannot get larger. I think I did not move the cavity, but I cannot find other reason for this situation.